With the estimates shown below, Sarah needs to determine the trade-in (replacement) value of machine X that will render its AW equal to that of machine Y at an interest rate of 9% per year. Determine the replacement value.

Machine X Machine Y
Market Value, $ ? 94,000
Annual Cost, $ per Year −57,500 −40,000 for year 1,increasing by 2000 per year thereafter.
Salvage Value 19,500 21,000
Life, Years 3 5

The replacement value is______ $ .
Answer #1

AW of machine Y = -94000 * (A/P, 9%,5) - 40000 - 2000 * (A/G, 9%,5) + 21000 * (A/F,9%,5)
= -94000 * 0.257092 - 40000 - 2000 * 1.828197 + 21000 * 0.167092

= -64314.14

Let trade in value be P, then

AW of machine X = - P *(A/P, 9%,3) - 57500 + 19500 * (A/F, 9%,3)

= - P * 0.395055 - 57500 + 19500 * 0.305055

= - P * 0.395055 - 51551.43

As per given condition

- P * 0.395055 - 51551.43 = -64314.14

P = (64314.14-51551.43)/0.395055 = 32306.16